OpenStudy (anonymous):

for the polynomial f(x)=-2x³-2x²+7x-25 as x→∞,f(x)→∞ a. true b. false

OpenStudy (anonymous):

\( \lim \limits_{x \to \infty} (-2 x^3 - 2 x^2 + 7 x - 25) = -\infty \)

OpenStudy (anonymous):

False. As x approaches infinity, \[-2x^{3}-2x^{2}\] will approach -infinity more rapidly than 7x approaches infinity, so \[f(x)\rightarrow \infty\] Basically, what FoolForMath said :P

OpenStudy (anonymous):

\[f(x)\rightarrow-\infty\] of course.